OpenStudy (anonymous):

the product of (3x + 7y)2

OpenStudy (anonymous):

\[\Large (3x+7y)*2\] ??

OpenStudy (anonymous):

two is multiplied or in exponent ???

OpenStudy (anonymous):

exponent

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

using formula \[\Large (a+b)^2=a^2+b^2+2ab\] here a=3x and b=7y \[\Large (3x+7y)^2=(3x)^2+(7y)^2+2(3x)(7y)\] or simplifying \[\Large =9x^2+49y^2+42xy\]

OpenStudy (anonymous):

@BrianaChantel did you get it ?

OpenStudy (anonymous):

Yea , Thanks

OpenStudy (anonymous):

you welcome :)